# Incentivizing Quality Text Generation via Statistical Contracts

### v1

@@ C.2 Multi-outcome contracts: Further Analysis
Table 5 shows the statistics of the various contract objectives in contracts when optimized without a monotonicity constraint, and displays how they match up to each other in cost-aware and cost-robust settings. We can observe that the min-pay contract minimizes expected pay at the expense of high budget and variance.The min-budget contract, on the other hand, is not the worst in any of the objectives. Additionally, the cost-robust setting sacrifices only a marginal increase in objective values: a $6.9\%$ increase in the Min-pay objective, an $8.7\%$ increase if optimizing for budget, and $6.5\%$ increase when optimizing for minimum variance.
@@ C.3 Implementation details
C.3.1 Contract design solvers

### v2

@@ E.2 Multi-outcome contracts: Further Analysis
Table 5 shows the statistics of the various contract objectives in contracts when optimized without a monotonicity constraint, and displays how they match up to each other in cost-aware and cost-robust settings. We observe that the min-pay contract minimizes expected pay at the expense of high budget and variance. The min-budget contract, on the other hand, is not the worst in any of the objectives. Additionally, the cost-robust setting sacrifices only a marginal increase in objective values: a $6.9\%$ increase in the Min-pay objective, an $8.7\%$ increase if optimizing for budget, and $6.5\%$ increase when optimizing for minimum variance.
@@ E.3 Implementation details
E.3.1 Contract design solvers